OpenStudy (anonymous):

solve by factoring. m^2 + 8m + 7 = 0

OpenStudy (anonymous):

m^2 + 8m + 7 = 0 -8 -8 m^2 + m + 7 = -8 -7 -7 m^2 +m = -15 m^3 = -15 understand?

OpenStudy (anonymous):

Hey sis

OpenStudy (anonymous):

Where'd that three come from? i understood up until that point & lulu shut up Lol

OpenStudy (anonymous):

Ohh , and thats not a option on here

OpenStudy (anonymous):

lol aha love u sis

OpenStudy (anonymous):

the 3 is the eponent and m^2 + m = m^3 because the it's like saying a.a.a how many a's? well three so m^3 are you still lost

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

can i get the optioin?

OpenStudy (anonymous):

8,7 -7,1 -7,-1 7,1

OpenStudy (anonymous):

oh ok so they want us to factor not solve ok sorry give me a sec to solve it

OpenStudy (anonymous):

it's 7 and 1 because when you write it like (7 + m)(1 + m)

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

Thanks , can you help with another ?

OpenStudy (anonymous):

Shabba Ranks

OpenStudy (anonymous):

bye @lulu22 Lmfao

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

One gold tooth

OpenStudy (anonymous):

n^2 + 2n - 24 = 0

OpenStudy (anonymous):

ok let me help now

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (anonymous):

ok go ahead

OpenStudy (anonymous):

n^2 + 2n - 24 = 0

OpenStudy (anonymous):

@Deivyneee

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (anonymous):

you want me to help?

OpenStudy (anonymous):

Yeah ,

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (anonymous):

ok give me time to solve

OpenStudy (anonymous):

Alright ,

OpenStudy (anonymous):

@Nora4Christ

OpenStudy (anonymous):

she still have not solve it

OpenStudy (anonymous):

A. m² + 8m + 7 = 0 Factor: (m + 7) (m + 1) = 0 Set each set equal to 0 m + 7 = 0 m + 1 = 0 And solve: m = -7 & m = -1

OpenStudy (anonymous):

sorry my connection was terrible @lulu22 @deivyneee